# Accelerated Convergence of Frank–Wolfe Algorithms with Adaptive Bregman Step-Size Strategy

### v1

@@ 6.4 Low-Rank Minimization
There exists a constant $L$ such that the pair $(f,\phi)$ is $L$-smad on $\mathbb{R}^{n}$ [Dragomir2021-rv]. Additionally, $f(X)$ is weakly convex on any compact set due to Proposition 2.10, which follows from the twice continuous differentiability of $f$.
We also compared BregFW with EucFW, ShortFW, and OpenFW. The parameter settings are the same as those in the previous subsection. We generated $X^{*}$ from an i.i.d. uniform distribution in $[0,1]$, normalized each column of $X^{*}$, and set $M=X^{*}(X^{*})^{\mathsf{T}}$. The initial point $X_{0}$ was generated from an i.i.d. uniform distribution in $[0,1]$. We set $\xi=10\lambda_{\max}(M)$ for $P$. Figure 5 shows the primal and FW gaps per iteration and gaps per second for $(n,r)=(1000,20)$ up to the 1000th iteration. Table 6 presents the average performance over 20 different instances for $(n,r)=(1000,20)$. BregFW performs slightly better than EucFW. OpenFW also performed as fast as BregFW and EucFW, but its performance was unstable. ShortFW did not converge due to the lack of Lipschitz continuity of $\nabla f$.
@@
Table 6: Average values of the primal gap, FW gap, and computational time (s) over 20 different instances of low-rank minimization for $(n,r)=(1000,20)$.
@@
BregFW 5.651947e-01 1.472666e+01 1.260218e+01
EucFW 5.776476e-01 1.723729e+01 1.228013e+01
ShortFW 2.710108e+08 2.169351e+09 4.137911e+00
OpenFW 1.978179e+00 2.121183e+03 4.176979e+00
6.5 Nonnegative Matrix Factorization
Given a nonnegative matrix $V\in\mathbb{R}_{+}^{m\times n}$, nonnegative matrix factorization (NMF) aims to find nonnegative matrices $W\in\mathbb{R}_{+}^{m\times r}$ and $H\in\mathbb{R}_{+}^{r\times n}$ such that $V\simeq WH$. NMF can be formulated as a minimization problem of the loss function that measures the difference between $V$ and $WH$, i.e.,
$\min_{(W,H)\in P}\quad f(W,H):=\frac{1}{2}\|WH-V\|_{F}^{2},$ (6.4)
where $P$ is a compact convex subset of $\mathbb{R}_{+}^{m\times r}\times\mathbb{R}_{+}^{r\times n}$. The objective function $f$ is weakly convex over $P$ due to Proposition 2.10. The gradient $\nabla f$ is not Lipschitz continuous, while $(f,\phi)$ is smooth adaptable [Mukkamala2019-mk] with $\phi(W,H)=\frac{1}{4}(\|W\|_{F}^{2}+\|H\|_{F}^{2})^{2}+\frac{1}{2}(\|W\|_{F}^{2}+\|H\|_{F}^{2})$.
@@ 6.5 Nonnegative Matrix Factorization
Figure 6: NMF for $(m,n,r)=(100,5000,20)$.
Table 7: Average values of the primal gap, FW gap, and computational time (s) over 20 different instances of NMF for $(m,n,r)=(100,5000,20)$.
@@
BregFW 1.201735e-04 1.151457e-02 5.098349e+01
EucFW 1.281873e-04 7.625941e-04 4.359247e+01
[…]
Figure 7: NMF for MovieLens 100K Dataset.
Acknowledgement
Appendix A Appendix
A.1 Results from Weak Convexity
Lemma A.1 (Primal gap bound from the quadratic growth).
$f(x)-f^{*}\leq\frac{2\mu}{(\mu-\rho)^{2}}\left(\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\right)^{2},$
or equivalently,
Proof.
$f(x)-f^{*}$ $=f(x)-f(x^{*})$
$\leq\langle\nabla f(x),x-x^{*}\rangle+\frac{\rho}{2}\|x-x^{*}\|^{2}$
which implies
$\left(1-\frac{\rho}{\mu}\right)\left(f(x)-f^{*}\right)\leq\langle\nabla f(x),x-x^{*}\rangle=\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\|x-x^{*}\|.$ (A.1)
Using the local quadratic growth condition again, we have
which provides
or equivalently
Lemma A.2 (Upper bound on primal gap for weakly convex functions).
Suppose that Assumptions 3.1 and 5.2 hold. Let $P$ be a polytope with the pyramidal width $\delta>0$. Let $\mathcal{S}$ denote any set of vertices of $P$ with $x\in\operatorname{conv}\mathcal{S}$. Let $\psi$ be any vector, so that we define $v^{\textup{FW}}=\operatorname*{argmin}_{v\in P}\langle\psi,v\rangle$ and $v^{\textup{A}}=\operatorname*{argmax}_{v\in\mathcal{S}}\langle\psi,v\rangle$. If $\rho/\mu<1$, then it holds that, for all $x\in[f\leq f^{*}+\zeta]\cap P$,
$f(x)-f^{*}\leq\frac{2\mu}{(\mu-\rho)^{2}\delta^{2}}\langle\nabla f(x),v^{\textup{A}}-v^{\textup{FW}}\rangle^{2}.$
Proof.
which implies
where the first inequality follows from $01-\rho/\mu>$ 0. Using (2.10) with $\psi=\nabla f(x)$ and $y=x^{*}$ and the above inequality, we obtain
$=\frac{\mu}{2}\left(1-\frac{\rho}{\mu}\right)^{2}(f(x)-f(x^{*})),$
A.2 Local Linear Convergence over Uniformly Convex Sets
To this end, we recall the definition of uniformly convex sets.
Definition A.3 ($(\alpha,p)$-uniformly convex set [Braun2022-ew, Definition 2.18][Kerdreux2021-kc, Definition 1.1]).
[…]
$y+\gamma(x-y)+\gamma(1-\gamma)\cdot\alpha\|x-y\|^{p}z\in P.$
Proposition A.4 (Scaling inequality [Braun2022-ew, Proposition 2.19][Kerdreux2021-kc, Lemma 2.1]).
$\frac{\langle u,x-v\rangle}{\|x-v\|^{p}}\geq\alpha\|u\|.$
Theorem A.5 (Local linear convergence over uniformly convex sets).
Suppose that Assumptions 3.1 and 5.2 with $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\operatorname{int}\operatorname{dom}\phi=\mathbb{R}^{n}$ hold and that $P$ is $(\alpha,p)$-uniformly convex set. Let $\nabla f$ be bounded away from 0, i.e., $\|\nabla f(x)\|\geq c>0$ for all $x\in P$. Let $D_{\textup{Euc}}:=\sup_{x,y\in P}\|x-y\|$ be the diameter of $P$. Consider the iterates of Algorithm 1 with $\gamma_{t}=\min\left\{\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L\|x_{t}-v_{t}\|^{2}},1\right\}$. Then, if $\rho<\mu$ and $\rho\leq L$, it holds that
for all $t\geq 1$ where
Proof.
Let $g_{t}:=\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle$ and $h_{t}:=f(x_{t})-f^{*}$. Lemma 3.3 with $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\nu=1$ is followed by $f(x_{t})-f(x_{t+1})\geq\frac{g_{t}}{2}\gamma_{t}$. Using Proposition A.4, we have
where the third inequality holds from Lemma 5.3 and $\|\nabla f(x)\|\geq c$, and the last inequality holds because of the local quadratic growth property of $f$. The initial bound $h_{1}\leq LD_{\textup{Euc}}^{2}=2LD^{2}$ holds from $\rho\leq L$ and (5.6). For $q=2$, we have $h_{t}-h_{t+1}\geq\frac{1}{2}\left(1-\frac{\rho}{\mu}\right)\min\left\{\frac{\alpha c}{L},1\right\}\cdot h_{t}$, which implies
Thus, we have the claim. For $p>2$, we use Lemma 2.20 with $c_{0}=LD_{\textup{Euc}}^{2}$, $c_{1}=\frac{1}{2}\left(1-\frac{\rho}{\mu}\right)$, $c_{2}=c_{1}\cdot\left(1-\frac{\rho}{\mu}\right)^{1-2/p}\frac{(\alpha c)^{2/p}}{L}$, and $\theta_{0}=1-2/p$ and obtain the claim. ∎
Note that Assumption 3.1 with $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\operatorname{int}\operatorname{dom}\phi=\mathbb{R}^{n}$ holds when $f$ is $L$-smooth over $P$. If $\rho\leq L$ does not hold, we can use the initial bound $h_{1}\leq\frac{\rho+L}{2}D_{\textup{Euc}}^{2}$ from (5.6). In that case, a local linear rate in Theorem A.5 is unchanged. Moreover, we have local linear convergence without assuming $\|\nabla f(x)\|>0$.
Theorem A.6 (Local linear convergence over uniformly convex sets without $\|\nabla f(x)\|>0$).
Suppose that Assumptions 3.1 and 5.2 with $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\operatorname{int}\operatorname{dom}\phi=\mathbb{R}^{n}$ hold and that $P$ is $(\alpha,p)$-uniformly convex set. Let $D_{\textup{Euc}}:=\sup_{x,y\in P}\|x-y\|$ be the diameter of $P$. Consider the iterates of Algorithm 1 with $\gamma_{t}=\min\left\{\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L\|x_{t}-v_{t}\|^{2}},1\right\}$. Then, if $\rho<\mu$ and $\rho\leq L$, it holds that
for all $t\geq 1$ where $c=\frac{\sqrt{2\mu}}{\mu-\rho}$ and
Proof.
Using an argument similar to Theorem A.5, we obtain
where the third inequality holds from the PL inequality (2.8) with $c=\frac{\sqrt{2\mu}}{\mu-\rho}$. Therefore, we use Lemma 2.20 with $c_{0}=LD_{\textup{Euc}}^{2}$, $c_{1}=\frac{1}{2}\left(1-\frac{\rho}{\mu}\right)$, $c_{2}=c_{1}\cdot\left(1-\frac{\rho}{\mu}\right)^{1-2/p}\frac{(\alpha c^{-1})^{2/p}}{L}$, and $\theta_{0}=1-1/p$ and obtain the claim. ∎

### v3

@@ E.4 Low-Rank Minimization
There exists a constant $L$ such that the pair $(f,\phi)$ is $L$-smad on $\mathbb{R}^{n}$ [Dragomir2021-rv]. Additionally, $f(X)$ is weakly convex on any compact set due to Proposition B.1, which follows from the twice continuous differentiability of $f$.
We also compared BregFW with EucFW, ShortFW, and OpenFW. The parameter settings are the same as those in the previous subsection. We generated $X^{*}$ from an i.i.d. uniform distribution in $[0,1]$, normalized each column of $X^{*}$, and set $M=X^{*}(X^{*})^{\mathsf{T}}$ (random seed 42). The initial point $X_{0}$ was generated from an i.i.d. uniform distribution in $[0,1]$. We set $\xi=10\lambda_{\max}(M)$ for $P$. Figure 8 shows the primal and FW gaps per iteration and gaps per second for $(n,r)=(1000,20)$ up to the 1000th iteration. Table 7 presents the average performance over 20 different instances for $(n,r)=(1000,20)$. BregFW performs slightly better than EucFW. OpenFW also performed as fast as BregFW and EucFW, but its performance was unstable. ShortFW did not converge due to the lack of Lipschitz continuity of $\nabla f$.
@@
Table 7: average numbers of the primal gap, FW gap, and computational time (s) over 20 different instances of low-rank minimization for $(n,r)=(1000,20)$. The results are reported as mean $\pm$ standard deviation.
@@
BregFW 5.6519e-01 $\pm$ 4.2e-02 1.4727e+01 $\pm$ 1.0e+00 1.2602e+01 $\pm$ 4.5e-01
EucFW 5.7765e-01 $\pm$ 4.6e-02 1.7237e+01 $\pm$ 2.0e+00 1.2280e+01 $\pm$ 3.8e-01
ShortFW 2.7101e+08 $\pm$ 1.9e+06 2.1694e+09 $\pm$ 1.5e+07 4.1379e+00 $\pm$ 7.9e-02
OpenFW 1.9782e+00 $\pm$ 3.6e-01 2.1212e+03 $\pm$ 3.1e+02 4.1770e+00 $\pm$ 1.7e-01
E.5 Nonnegative Matrix Factorization
Given a nonnegative matrix $V\in\mathbb{R}_{+}^{m\times n}$, nonnegative matrix factorization (NMF) aims to find nonnegative matrices $W\in\mathbb{R}_{+}^{m\times r}$ and $H\in\mathbb{R}_{+}^{r\times n}$ such that $V\simeq WH$. NMF can be formulated as a minimization problem of the loss function that measures the difference between $V$ and $WH$, i.e., $\min_{(W,H)\in P}f(W,H):=\frac{1}{2}\|WH-V\|_{F}^{2}$, where $P$ is a compact convex subset of $\mathbb{R}_{+}^{m\times r}\times\mathbb{R}_{+}^{r\times n}$. The objective function $f$ is weakly convex over $P$ due to Proposition B.1. The gradient $\nabla f$ is not Lipschitz continuous, while $(f,\phi)$ is smooth adaptable [Mukkamala2019-mk] with $\phi(W,H)=\frac{1}{4}(\|W\|_{F}^{2}+\|H\|_{F}^{2})^{2}+\frac{1}{2}(\|W\|_{F}^{2}+\|H\|_{F}^{2})$.
@@ E.5 Nonnegative Matrix Factorization
[…]
Figure 9: Log-log plot of the primal and the FW gaps on NMF for $(m,n,r)=(100,5000,20)$.
Figure 10: Log-log plot of the primal and the FW gaps on NMF for MovieLens 100K Dataset.
Table 8: average numbers of the primal gap, FW gap, and computational time (s) over 20 different instances of NMF for $(m,n,r)=(100,5000,20)$. The results are reported as mean $\pm$ standard deviation.
@@
BregFW 1.2017e-04 $\pm$ 5.7e-06 1.1515e-02 $\pm$ 1.4e-02 5.0983e+01 $\pm$ 1.1e+00
EucFW 1.2819e-04 $\pm$ 4.6e-06 7.6259e-04 $\pm$ 8.0e-04 4.3592e+01 $\pm$ 1.0e+00